# The possible adelic indices for elliptic curves admitting a rational cyclic isogeny

### v1

@@ 3.1. Map to a Rank 0 Elliptic Curve
Sometimes a modular curve $X$ admits a nonconstant rational map to an elliptic curve $E$ of rank 0. In this case, we use the code of the third-named author and Rouse [JacobJeremycode] to compute the rational points of $X$. The function EllipticCurveQuoCandidates lists candidates for elliptic quotients of the Jacobian of $X$. The function FindMapsToEC then computes an explicit map from $X$ to an elliptic curve $E^{\prime}$ that is isogenous to one of these candidates. Broadly speaking, the code works by selecting a weight $2$ cusp form $f$ whose Hecke eigenvalues agree with those of $E$, computing period integrals of $f$ numerically to recover its period lattice up to a rational scalar, and then using linear algebra on Fourier expansions to express the coordinate functions $x$ and $y$ as ratios of modular forms for $X$, thereby obtaining a map $X\to E^{\prime}$. After a suitable translation, we obtain a map defined over $\mathbb{Q}$. Although the intermediate steps in this function are heuristic, the final map is verified algebraically and is therefore rigorous. Once a map is obtained, we call the function RatPtsFromMaps to determine the rational points of $X$ by pulling back the finitely many rational points of $E^{\prime}$ using Hensel lifting and rational reconstruction.
@@ 3.2. Sieving with Respect to a Projection
This method is based on Section 8.3 of [MR4468989]. For some modular curves $X$, we can find an automorphism $i\colon X\to X$ such that the Jacobian $J_{X}$ of $X$ and the Jacobian of the quotient curve have the same rank. When this occurs, there exists an abelian variety $V$ of rank 0 such that $J_{X}$ is the product of $V$ and the Jacobian of the quotient curve. For any point $P\in X(\mathbb{Q})$, the point Pi(P)P-i(P) lies in $J_{X}(\mathbb{Q})_{\operatorname{tors}}$. Hence, it suffices to compute the preimages of rational points in $J_{X}(\mathbb{Q})_{\operatorname{tors}}$ under the map $a\colon X(\mathbb{Q})\to J_{X}(\mathbb{Q})_{\operatorname{tors}}$ given by $P\mapsto P-i(P)$, which is injective away from the fixed points of $i$.
In practice, it is difficult to determine $J_{X}(\mathbb{Q})_{\operatorname{tors}}$. However, by the appendix of Katz’s article [MR604840], if $p$ is an odd prime of good reduction, then $J_{X}(\mathbb{Q})_{\operatorname{tors}}$ injects into $J_{X}(\mathbb{F}_{p})$. We can use this to obtain an upper bound on the order of $J_{X}(\mathbb{Q})_{\operatorname{tors}}$. After this, we either find a point $P\in X(\mathbb{Q})$ such that Pi(P)P-i(P) has a given order or we find a prime $p$ and show that, modulo $p$, there is no divisor of the form PQP-Q of that order for any $P,Q\in X(\mathbb{F}_{p})$.
@@ 3.3. The Method of Chabauty and Coleman
This is a method for determining the rational points on a curve $X$ when the rank of its Jacobian is less than its genus. When this assumption holds, one can define a finite subset of the Jacobian of $X$ over $\mathbb{Q}_{p}$ that contains the set $X(\mathbb{Q})$. This method is originally due to Chabauty [chabauty1941points] and was made effective by Coleman [coleman1985effective]. We will apply this to several genus 2 curves. There are two built-in commands in Magma that implement this method.
@@ 3.4. Quadratic Chabauty and the Mordell–Weil Sieve
We apply quadratic Chabauty along with the Mordell–Weil sieve to determine the rational points of certain modular curves where the rank of the Jacobian is equal to the genus of the curve. Analogously to the creation of a subset of $J(\mathbb{Q}_{p})$ in the case of classical Chabauty–Coleman, quadratic Chabauty uses $p$-adic heights to construct a finite subset of $p$-adic points that contains the set of rational points. For more details, see [MR2512779] and [MR3843370]. Recent work [bianchi2022rational, finnerty2025quadraticchabautyexperimentsgenus] has specifically addressed the case of bielliptic modular curves of genus $2$ and rank 2, which arises on a few occasions in this paper.
The Mordell–Weil sieve was originally used in Scharaschkin’s Ph.D. thesis [scharaschkinthesis] to prove that a curve has no rational points, but it can be used more generally to determine the set of rational points. In particular, it can eliminate $p$-adic points from the output of classical or quadratic Chabauty by proving that they do not correspond to rational points.
@@
We know points in $A_{i}$ only modulo $p_{i}^{m_{i}}$ for some integer $m_{i}$ up to a chosen precision, and we can compute $a_{1}$ and $a_{2}$ modulo $p_{i}^{n_{i}}$ for some $n_{i}\leq m_{i}$. This means that for every point in $A_{i}$, there are at most $|J(\mathbb{Q})_{\operatorname{tors}}|$ possibilities for the image of $\iota(P)$ in $J(\mathbb{Q})/p_{i}^{n_{i}}J(\mathbb{Q})$. Our goal in the Mordell–Weil sieve is to show that these cosets in $J(\mathbb{Q})/p_{i}^{n_{i}}J(\mathbb{Q})$ cannot arise from any rational point by considering reduction maps modulo multiple primes. We follow the approach of Bianchi and Padurariu [bianchi2022rational].
@@
This may require trying both different eligible sets $S$ and auxiliary integers $M^{\prime}$. We carry out this computation explicitly, with documentation available at [finnertysievecode]. In each case to which we applied this method, the sieve eliminated all points that did not easily lift to rational points. Table 1 shows the resulting affine rational points on each of the curves used in this analysis.
@@ 4.1. Strategy and Preliminary Results
Lemos’s theorem (Theorem 3) gives that the $\ell$-adic representation of $E$ is surjective for all primes $\ell>37$, so we restrict our attention to primes $\ell\leq 37$. By Theorem 1.1.6 and Table 1 of [MR4468989], for each prime $p\in\{2,3,5,7,13\}$, the possible $\ell$-adic images of $E$ (with $\ell\leq 37$) arise as non-CM rational points on the fiber products $X_{0}(p)$ with the modular curves in the following table:
@@
Several of the modular curves in the table above have no known non-CM rational points. In fact, it is conjectured that for $\ell\geq 19$, the curve $X_{\mathrm{ns}}^{+}(\ell)$ has no non-CM rational points (see, for instance, [zywina2015possibleimagesmodell, Conjecture 1.12]). However, to obtain an unconditional result, we must still consider fiber products with these modular curves.
@@
This means that for any positive integer $m$, we can always consider a twist $E^{\prime}$ of $E$ with $I\in G_{E^{\prime}}(m)$ by Lemma 11. This greatly reduces the number of possible $\ell$-adic images to consider, since we need only consider those containing I-I. We also use the following result of Lemos [MR3885140] throughout this section when considering fiber products involving $X_{0}(p)$.
@@
The argument in [MR3885140, Section 3] shows that the $j$-invariant of $E$ belongs to $\mathbb{Z}[\frac{1}{\ell}]$ as long as $\ell\neq p$. This pair $(E,\ell)$ also satisfies the conditions of [MR3885140, Proposition 2.2], implying that $E$ has potentially good reduction at $\ell$. Hence $j(E)\in\mathbb{Z}[\frac{1}{\ell}]\cap\mathbb{Z}_{\ell}=\mathbb{Z}$, which completes the proof. ∎
We now prove two lemmas that will help us rule out non-CM rational points on fiber products with the modular curves $X_{\mathrm{ns}}^{+}(\ell)$ and $X_{S_{4}}(13)$. All of the computational claims made in this section can be verified in the file Section 4.1.m in the repository [projectcode].
@@
Let $p\in\{2,3,5,7,13\}$ and $\ell>5$ be a prime distinct from $p$. From Proposition 16, we know that if $X_{0}(p)\times X_{\mathrm{ns}}^{+}(\ell)$ has any non-CM rational points, then the corresponding $j$-invariant must be an integer. The set of integral $j$-invariants that occur for non-CM rational points of $X_{0}(p)$ is finite for each $p$ and listed in [MR3885140, p. 7]. We reproduce these $j$-invariants in the table below.
@@
From Section 5.1 of [MR4589060], we know that the modular curve $X_{S_{4}}(13)$ (13.91.3.a.1) has three non-CM $j$-invariants:
@@
For each of these $j$-invariants, we consider a representative elliptic curve $E/\mathbb{Q}$. The LMFDB labels of the first two representatives are 50700.z1 and 61347.bb1, respectively. Let $p\neq 13$ be a prime. For both of these curves, we see that their $p$-adic representation is surjective. Hence, if any elliptic curve $E/\mathbb{Q}$ has either of these two $j$-invariants, then its $p$-adic representation must also be surjective (this follows by [GMR_smallestSurjectivePrime, Lemma 5]). The result follows similarly for the third $j$-invariant; however, there is no elliptic curve in the LMFDB with this $j$-invariant, so we instead construct one using the Magma command EllipticCurveFromjInvariant and run Zywina’s FindOpenImage [ZywinaOpen] to compute its Galois image. We find that the only prime divisors of its adelic level are among $2$, 55, $13$, $61$ and $109$. If $p$ is any of these primes (except $13$), we check that the $p$-adic representation is surjective, and if $p$ is outside this list, then necessarily the $p$-adic representation is surjective. ∎
@@ 4.2. Computation of Possible Images
Part (1) follows from Tables 11-14 of [MR4868206]. Part (2) follows from Theorem 1.1 of [rakvi2023classification3adicgalois]. Part (3) follows for primes $\ell>37$ by Theorem 3, so it remains to treat the primes $5\leq\ell\leq 37$. We proceed along the lines of Table 2, using the strategy set forth in Section 4.1.
@@
By our assumption that $E$ does not admit a rational isogeny of prime degree greater than $2$, we may disregard the fiber product $X_{0}(2)\times X_{0}(5)$. The curves $X_{0}(2)\times X_{S_{4}}(5)$ and $X_{0}(2)\times X_{\mathrm{ns}}^{+}(5)$ have LMFDB labels 10.15.1.a.1 and 10.30.1.a.1, respectively. They both have genus 1 and rank 0. We check that each has no non-CM rational points. It follows that the $5$-adic Galois representation of $E$ is surjective.
@@
Again, by our isogeny assumption, we do not need to consider the fiber product $X_{0}(2)\times X_{0}(7)$. The curve $X_{0}(2)\times X_{\mathrm{ns}}^{+}(7)$ (14.63.2.a.1) has no non-CM rational points by Lemma 17. The curve $X_{0}(2)\times X_{\mathrm{sp}}^{+}(7)$ (14.84.3.a.1) has genus $3$ and rank 0. It admits a map to a rank 0 elliptic curve. Running the code ModCrvToEC [JacobJeremycode] (discussed in Section 3.1), we determine that this curve has exactly four rational points, all of which are cusps or CM points. Hence the $7$-adic Galois representation of $E$ is surjective.
@@
All computations can be verified in the file 2 isogeny l adic images.m in the repository [projectcode]. ∎
@@
    If the $2$-adic Galois image of $E$ contains I-I, then it is conjugate to one of the following subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$: $\operatorname{GL}_{2}(\mathbb{Z}_{2})$, 2.2.0.a.1, 4.2.0.a.1, 8.2.0.a.1, 8.2.0.b.1, 2.3.0.a.1, 2.6.0.a.1, 4.6.0.b.1, 4.6.0.a.1, 4.6.0.c.1, 8.6.0.d.1, 8.6.0.c.1, 8.6.0.b.1,
    8.6.0.a.1, or 4.8.0.b.1.
@@
    If the $3$-adic Galois image of $E$ contains I-I, then it is conjugate to one of the following subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{3})$: 3.4.0.a.1, 3.12.0.a.1, 9.12.0.a.1, 9.12.0.b.1, 9.36.0.a.1, 9.36.0.b.1, 9.36.0.c.1, 9.36.0.d.1, 9.36.0.d.2, 9.36.0.e.1, 9.36.0.f.2,
    9.36.0.g.1, 9.36.0.f.1, or 27.36.0.a.1.
@@
Part (1) follows from Tables 15-19 of [MR4868206]. Part (2) follows from Theorem 1.1 of [rakvi2023classification3adicgalois]. For primes $\ell>37$, part (4) follows from Theorem 1.1 of [MR3885140].
@@
All computations can be verified in the file 3 isogeny l adic images.m in the repository [projectcode]. ∎
@@
    If the $2$-adic Galois image of $E$ contains I-I, then it is conjugate to one of the following subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$: $\operatorname{GL}_{2}(\mathbb{Z}_{2})$, 2.3.0.a.1, 4.4.0.a.1, 8.2.0.a.1, 8.6.0.f.1, 8.6.0.d.1, or 8.6.0.a.1.
@@
Part (1) follows from Tables 17-19 of [MR4868206]. Part (2) follows from Theorem 1.1 of [rakvi2023classification3adicgalois]. Part (3) follows from Theorem 1.1.6 of [MR4468989]. Part (4) follows for primes $\ell>37$ by Theorem 3.
Proving $7$-adic surjectivity. The fiber product $X_{0}(5)\times X_{0}(7)$ has label 35.48.3.a.1; it has no non-CM rational points by Theorem 6. Next, we discuss the curve $X_{0}(5)\times X_{\mathrm{ns}}^{+}(7)$ (35.126.6.a.1). Let $C$ denote the subscheme of $\mathbb{P}^{3}_{\mathbb{Q}}$ defined by the sequence of homogeneous polynomials giving the canonical model of 35.126.6.a.1. We verify that the set $C(\mathbb{Z}/49\mathbb{Z})$ is empty; hence, there are no rational points on $X_{0}(5)\times X_{\mathrm{ns}}^{+}(7)$. Lastly, seeking a contradiction, suppose that $E/\mathbb{Q}$ corresponds to a rational point on the fiber product $X_{0}(5)\times X_{\mathrm{sp}}^{+}(7)$ (35.168.9.a.1). Then there exists a quadratic field $K$ such that $E/K$ corresponds to a quadratic point on the fiber product $X_{0}(5)\times X_{\mathrm{sp}}(7)$. In particular, $E/K$ corresponds to a quadratic point on the fiber product $X_{0}(5)\times X_{0}(7)$. Moreover, since $E$ is defined over $\mathbb{Q}$, its $j$-invariant lies in $\mathbb{Q}$. But by Theorem $1$ of [Vux91], this is impossible.
@@
All computations can be verified in the file 5 isogeny l adic images.m in the repository [projectcode]. ∎
@@
Part (1) follows from Table 17 and Table 19 of [MR4868206]. Part (2) follows from Theorem 1.1 of [rakvi2023classification3adicgalois]. Part (3) follows from Theorem 1.1.6 of [MR4468989]. For primes $\ell>37$, part (4) follows from Theorem 1.1 of [MR3885140]. Thus, it only remains to verify part (4) for primes $\ell\leq 37$.
@@
Let us next consider the fiber product $X_{0}(7)\times X_{\mathrm{ns}}^{+}(5)$ (35.80.5.a.1). This curve, which we denote by $C$, has genus 5 and rank 1. We use the canonical model given on LMFDB. It has two known rational points $P_{1}=(1,-6,3,3,5)$ and $P_{2}=(1,-2,3,7,-5)$; these both correspond to CM elliptic curves. We will now prove that these are the only rational points. There exists an automorphism $i\colon C(\mathbb{Q})\to C(\mathbb{Q})$ given by
$x,y,z,w,t]\mapsto[x,y,z,w,x-2y+2z-3w-t].$
@@
All computations can be verified in the file 7 isogeny l adic images.m in the repository [projectcode]. ∎
@@
    If the $13$-adic Galois image of $E$ contains I-I, then it is conjugate to one of the following subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{13})$: 13.14.0.a.1, 13.28.0.a.1, 13.28.0.a.2, 13.42.0.a.2, 13.42.0.a.1, or 13.42.0.b.1.
@@
Part (1) follows from Table 19 of [MR4868206]. Part (2) follows from Theorem 1.1.6 of [MR4468989]. In the case that $\ell=3$, part (3) follows from Theorem 1.1 of [rakvi2023classification3adicgalois]. For primes $\ell>37$, part (3) follows from Theorem 1.1 of [MR3885140]. Thus, it only remains to verify part (3) for primes $\ell\leq 37$.

### v2

@@ 3.1. Map to a Rank 0 Elliptic Curve
Sometimes a modular curve $X$ admits a nonconstant rational map to an elliptic curve $E$ defined over $\mathbb{Q}$ of rank 0. In this case, we use the code of the third-named author and Rouse [31, 32] to find such a map and compute the rational points of $X$. We first compute a model of $X$ and Fourier expansions of weight $2$ cusp forms using Zywina’s code [47]. We attempt to fix a rational base point $Q\in X(\mathbb{Q})$ (if none is found, we can attempt to proceed using the cusp at infinity, which need not be rational).
We identify a weight $2$ cusp form $f$ whose Hecke eigenvalues for primes $p\nmid N$ agree with the corresponding Fourier coefficients $a_{p}(E)$. We then compute periods of $f$ and (up to a scalar) match the resulting period lattice with the period lattice $\Lambda(E^{\prime})$ of an optimal elliptic curve $E^{\prime}$ in the isogeny class of $E$. We consider the natural identification
$\alpha\colon X(\mathbb{C})\longrightarrow\mathbb{H}/(G\cap\operatorname{SL}_{2}(\mathbb{Z})),$
where $G$ is the subgroup associated to the modular curve $X$. Using the chosen base point, we obtain an analytic map $X\to\mathbb{C}/\Lambda(E^{\prime})$ given by
$P\longmapsto\int_{\alpha(Q)}^{\alpha(P)}f\,ds,$
and composing with the natural isomorphism $\mathbb{C}/\Lambda(E^{\prime})\simeq E^{\prime}(\mathbb{C})$ gives a map $X\to E^{\prime}$ defined over a cyclotomic field. When a rational base point $Q\in X(\mathbb{Q})$ is available, we translate so that $Q$ maps to the origin and then recover explicit defining equations for a map $X\to E^{\prime}$ defined over $\mathbb{Q}$.
Although the construction uses numerical and heuristic steps, we then rigorously verify that the resulting polynomials indeed define a nonconstant map $X\to E^{\prime}$ as claimed. Once the map is verified, we determine $X(\mathbb{Q})$ by pulling back each rational point $T\in E^{\prime}(\mathbb{Q})$ and determining the rational points on the resulting zero dimensional fibers. When possible, the method described in Step 13 of [32] (based on $p$-adic lifting together with point counting bounds) is used to do this efficiently. Otherwise, we use Magma’s built-in commands for determining the rational points of the zero dimensional fibers, which work well when the genus of $X$ is low but are slow in higher genus.
@@ 3.2. Sieving with Respect to a Projection
This method is based on Section 8.3 of [36]. For some modular curves $X$, we can find an automorphism $i\colon X\to X$ such that the Jacobian $J_{X}$ of $X$ and the Jacobian of the quotient curve have the same rank. When this occurs, there exists an abelian variety $V$ of rank 0 such that $J_{X}$ is the product of $V$ and the Jacobian of the quotient curve. For any point $P\in X(\mathbb{Q})$, the point Pi(P)P-i(P) lies in $J_{X}(\mathbb{Q})_{\operatorname{tors}}$. Hence, it suffices to compute the preimages of rational points in $J_{X}(\mathbb{Q})_{\operatorname{tors}}$ under the map $a\colon X(\mathbb{Q})\to J_{X}(\mathbb{Q})_{\operatorname{tors}}$ given by $P\mapsto P-i(P)$, which is injective away from the fixed points of $i$.
In practice, it is difficult to determine $J_{X}(\mathbb{Q})_{\operatorname{tors}}$. However, by the appendix of Katz’s article [24], if $p$ is an odd prime of good reduction, then $J_{X}(\mathbb{Q})_{\operatorname{tors}}$ injects into $J_{X}(\mathbb{F}_{p})$. We can use this to obtain an upper bound on the order of $J_{X}(\mathbb{Q})_{\operatorname{tors}}$. After this, we either find a point $P\in X(\mathbb{Q})$ such that Pi(P)P-i(P) has a given order or we find a prime $p$ and show that, modulo $p$, there is no divisor of the form PQP-Q of that order for any $P,Q\in X(\mathbb{F}_{p})$.
@@ 3.3. The Method of Chabauty and Coleman
This is a method for determining the rational points on a curve $X$ when the rank of its Jacobian is less than its genus. When this assumption holds, one can define a finite subset of the Jacobian of $X$ over $\mathbb{Q}_{p}$ that contains the set $X(\mathbb{Q})$. This method is originally due to Chabauty [11] and was made effective by Coleman [14]. We will apply this to several genus 2 curves. There are two built-in commands in Magma that implement this method.
@@ 3.4. Quadratic Chabauty and the Mordell–Weil Sieve
We apply quadratic Chabauty along with the Mordell–Weil sieve to determine the rational points of certain modular curves where the rank of the Jacobian is equal to the genus of the curve. Analogously to the creation of a subset of $J(\mathbb{Q}_{p})$ in the case of classical Chabauty–Coleman, quadratic Chabauty uses $p$-adic heights to construct a finite subset of $p$-adic points that contains the set of rational points. For more details, see [26] and [3]. Recent work [5, 18] has specifically addressed the case of bielliptic modular curves of genus $2$ and rank 2, which arises on a few occasions in this paper.
The Mordell–Weil sieve was originally used in Scharaschkin’s Ph.D. thesis [38] to prove that a curve has no rational points, but it can be used more generally to determine the set of rational points. In particular, it can eliminate $p$-adic points from the output of classical or quadratic Chabauty by proving that they do not correspond to rational points.
@@
We know points in $A_{i}$ only modulo $p_{i}^{m_{i}}$ for some integer $m_{i}$ up to a chosen precision, and we can compute $a_{1}$ and $a_{2}$ modulo $p_{i}^{n_{i}}$ for some $n_{i}\leq m_{i}$. This means that for every point in $A_{i}$, there are at most $|J(\mathbb{Q})_{\operatorname{tors}}|$ possibilities for the image of $\iota(P)$ in $J(\mathbb{Q})/p_{i}^{n_{i}}J(\mathbb{Q})$. Our goal in the Mordell–Weil sieve is to show that these cosets in $J(\mathbb{Q})/p_{i}^{n_{i}}J(\mathbb{Q})$ cannot arise from any rational point by considering reduction maps modulo multiple primes. We follow the approach of Bianchi and Padurariu [5].
@@
This may require trying both different eligible sets $S$ and auxiliary integers $M^{\prime}$. We carry out this computation explicitly, with documentation available at [17]. In each case to which we applied this method, the sieve eliminated all points that did not easily lift to rational points. Table 1 shows the resulting affine rational points on each of the curves used in this analysis.
@@ 4.1. Strategy and Preliminary Results
Lemos’s theorem (Theorem 3) gives that the $\ell$-adic representation of $E$ is surjective for all primes $\ell>37$, so we restrict our attention to primes $\ell\leq 37$. By Theorem 1.1.6 and Table 1 of [36], for each prime $p\in\{2,3,5,7,13\}$, the possible $\ell$-adic images of $E$ (with $\ell\leq 37$) arise as non-CM rational points on the fiber products $X_{0}(p)$ with the modular curves in the following table:
@@
Several of the modular curves in the table above have no known non-CM rational points. In fact, it is conjectured that for $\ell\geq 19$, the curve $X_{\mathrm{ns}}^{+}(\ell)$ has no non-CM rational points (see [44, Conjecture 1.12], for instance). However, to obtain an unconditional result, we must still consider fiber products with these modular curves.
@@
This means that for any positive integer $m$, we can always consider a twist $E^{\prime}$ of $E$ with $I\in G_{E^{\prime}}(m)$ by Lemma 10. This greatly reduces the number of possible $\ell$-adic images to consider, since we need only consider those containing I-I. We also use the following result of Lemos [28] throughout this section when considering fiber products involving $X_{0}(p)$.
@@
The argument in [28, Section 3] shows that the $j$-invariant of $E$ belongs to $\mathbb{Z}[\frac{1}{\ell}]$ as long as $\ell\neq p$. This pair $(E,\ell)$ also satisfies the conditions of [28, Proposition 2.2], implying that $E$ has potentially good reduction at $\ell$. Hence $j(E)\in\mathbb{Z}[\frac{1}{\ell}]\cap\mathbb{Z}_{\ell}=\mathbb{Z}$, which completes the proof. ∎
We now prove two lemmas that will help us rule out non-CM rational points on fiber products with the modular curves $X_{\mathrm{ns}}^{+}(\ell)$ and $X_{S_{4}}(13)$. All of the computational claims made in this section can be verified in the file Section 4.1.m in the repository [16].
@@
Let $p\in\{2,3,5,7,13\}$ and $\ell>5$ be a prime distinct from $p$. From Proposition 15, we know that if $X_{0}(p)\times X_{\mathrm{ns}}^{+}(\ell)$ has any non-CM rational points, then the corresponding $j$-invariant must be an integer. The set of integral $j$-invariants that occur for non-CM rational points of $X_{0}(p)$ is finite for each $p$ and listed in [28, p. 7]. We reproduce these $j$-invariants in the table below.
@@
From Section 5.1 of [2], we know that the modular curve $X_{S_{4}}(13)$ (13.91.3.a.1) has three non-CM $j$-invariants:
@@
For each of these $j$-invariants, we consider a representative elliptic curve $E/\mathbb{Q}$. The LMFDB labels of the first two representatives are 50700.z1 and 61347.bb1, respectively. Let $p\neq 13$ be a prime. For both of these curves, we see that their $p$-adic representation is surjective. Hence, if any elliptic curve $E/\mathbb{Q}$ has either of these two $j$-invariants, then its $p$-adic representation must also be surjective (this follows by [20, Lemma 5]). The result follows similarly for the third $j$-invariant; however, there is no elliptic curve in the LMFDB with this $j$-invariant, so we instead construct one using the Magma command EllipticCurveFromjInvariant and run Zywina’s FindOpenImage [46] to compute its Galois image. We find that the only prime divisors of its adelic level are among $2$, 55, $13$, $61$ and $109$. If $p$ is any of these primes (except $13$), we check that the $p$-adic representation is surjective, and if $p$ is outside this list, then necessarily the $p$-adic representation is surjective. ∎
@@ 4.2. Computation of Possible Images
Part (1) follows from Tables 11-14 of [13]. Part (2) follows from Theorem 1.1 of [35]. Part (3) follows for primes $\ell>37$ by Theorem 3, so it remains to treat the primes $5\leq\ell\leq 37$. We proceed along the lines of Table 2, using the strategy set forth in Section 4.1.
@@
By our assumption that $E$ does not admit a rational isogeny of prime degree greater than 22, we may disregard the fiber product $X_{0}(2)\times X_{0}(5)$. The curves $X_{0}(2)\times X_{S_{4}}(5)$ and $X_{0}(2)\times X_{\mathrm{ns}}^{+}(5)$ have LMFDB labels 10.15.1.a.1 and 10.30.1.a.1, respectively. They both have genus 1 and rank 0. We check that each has no non-CM rational points. It follows that the $5$-adic Galois representation of $E$ is surjective.
@@
Again, by our isogeny assumption, we do not need to consider the fiber product $X_{0}(2)\times X_{0}(7)$. The curve $X_{0}(2)\times X_{\mathrm{ns}}^{+}(7)$ (14.63.2.a.1) has no non-CM rational points by Lemma 16. The curve $X_{0}(2)\times X_{\mathrm{sp}}^{+}(7)$ (14.84.3.a.1) has genus $3$ and rank 0. It admits a map to a rank 0 elliptic curve. Running the code ModCrvToEC [31] (discussed in Section 3.1), we determine that this curve has exactly four rational points, all of which are cusps or CM points. Hence the $7$-adic Galois representation of $E$ is surjective.
@@
All computations can be verified in the file 2 isogeny l adic images.m in the repository [16]. ∎
@@
    If the $2$-adic Galois image of $E$ contains I-I, then it is conjugate to one of the following subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$: $\operatorname{GL}_{2}(\mathbb{Z}_{2})$, 2.2.0.a.1, 2.3.0.a.1, 2.6.0.a.1, 4.2.0.a.1, 4.6.0.a.1, 4.6.0.b.1, 4.6.0.c.1, 4.8.0.b.1, 8.2.0.a.1, 8.2.0.b.1, 8.6.0.a.1, 8.6.0.b.1,
    8.6.0.c.1, or 8.6.0.d.1.
@@
    If the $3$-adic Galois image of $E$ contains I-I, then it is conjugate to one of the following subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{3})$: 3.4.0.a.1, 3.12.0.a.1, 9.12.0.a.1, 9.12.0.b.1, 9.36.0.a.1, 9.36.0.b.1, 9.36.0.c.1, 9.36.0.d.1, 9.36.0.d.2, 9.36.0.e.1, 9.36.0.f.1,
    9.36.0.f.2, 9.36.0.g.1, or 27.36.0.a.1.
@@
Part (1) follows from Tables 15-19 of [13]. Part (2) follows from Theorem 1.1 of [35]. For primes $\ell>37$, part (4) follows from Theorem 1.1 of [28].
@@
All computations can be verified in the file 3 isogeny l adic images.m in the repository [16]. ∎
@@
    If the $2$-adic Galois image of $E$ contains I-I, then it is conjugate to one of the following subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{2})$: $\operatorname{GL}_{2}(\mathbb{Z}_{2})$, 2.3.0.a.1, 4.4.0.a.1, 8.2.0.a.1, 8.6.0.a.1, 8.6.0.d.1, or 8.6.0.f.1.
@@
Part (1) follows from Tables 17-19 of [13]. Part (2) follows from Theorem 1.1 of [35]. Part (3) follows from Theorem 1.1.6 of [36]. Part (4) follows for primes $\ell>37$ by Theorem 3.
Proving $7$-adic surjectivity. The fiber product $X_{0}(5)\times X_{0}(7)$ has label 35.48.3.a.1; it has no non-CM rational points by Theorem 5. Next, we discuss the curve $X_{0}(5)\times X_{\mathrm{ns}}^{+}(7)$ (35.126.6.a.1). Let $C$ denote the subscheme of $\mathbb{P}^{3}_{\mathbb{Q}}$ defined by the sequence of homogeneous polynomials giving the canonical model of 35.126.6.a.1. We verify that the set $C(\mathbb{Z}/49\mathbb{Z})$ is empty; hence, there are no rational points on $X_{0}(5)\times X_{\mathrm{ns}}^{+}(7)$. Lastly, seeking a contradiction, suppose that $E/\mathbb{Q}$ corresponds to a rational point on the fiber product $X_{0}(5)\times X_{\mathrm{sp}}^{+}(7)$ (35.168.9.a.1). Then there exists a quadratic field $K$ such that $E/K$ corresponds to a quadratic point on the fiber product $X_{0}(5)\times X_{\mathrm{sp}}(7)$. In particular, $E/K$ corresponds to a quadratic point on the fiber product $X_{0}(5)\times X_{0}(7)$. Moreover, since $E$ is defined over $\mathbb{Q}$, its $j$-invariant lies in $\mathbb{Q}$. But by Theorem $1$ of [43], this is impossible.
@@
All computations can be verified in the file 5 isogeny l adic images.m in the repository [16]. ∎
@@
Part (1) follows from Table 17 and Table 19 of [13]. Part (2) follows from Theorem 1.1 of [35]. Part (3) follows from Theorem 1.1.6 of [36]. For primes $\ell>37$, part (4) follows from Theorem 1.1 of [28]. Thus, it only remains to verify part (4) for primes $\ell\leq 37$.
@@
Let us next consider the fiber product $X_{0}(7)\times X_{\mathrm{ns}}^{+}(5)$ (35.80.5.a.1). This curve, which we denote by $C$, has genus 5 and rank 1. We use the canonical model in $\mathbb{P}^{4}$ given on LMFDB. It has two known rational points $P_{1}=(1:-6:3:3:5)$ and $P_{2}=(1:-2:3:7:-5)$; these both correspond to CM elliptic curves. We will now prove that these are the only rational points. There exists an automorphism $i\colon C(\mathbb{Q})\to C(\mathbb{Q})$ given by
$(x:y:z:w:t)\mapsto(x:y:z:w:x-2y+2z-3w-t).$
@@
All computations can be verified in the file 7 isogeny l adic images.m in the repository [16]. ∎
@@
    If the $13$-adic Galois image of $E$ contains I-I, then it is conjugate to one of the following subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{13})$: 13.14.0.a.1, 13.28.0.a.1, 13.28.0.a.2, 13.42.0.a.1, 13.42.0.a.2, or 13.42.0.b.1.
@@
Part (1) follows from Table 19 of [13]. Part (2) follows from Theorem 1.1.6 of [36]. In the case that $\ell=3$, part (3) follows from Theorem 1.1 of [35]. For primes $\ell>37$, part (3) follows from Theorem 1.1 of [28]. Thus, it only remains to verify part (3) for primes $\ell\leq 37$.